The bulk modulus of water is 1.9GPa. Compute the volume contraction of 65 mL of water when subjected to a pressure of 3.3 MPa.


1
Expert's answer
2021-07-20T14:58:19-0400

We know that

bulk modulus

"\beta=\frac{∆P}{\frac{∆V}{V}}\rightarrow(1)"

"∆P=3.3MPa\\V=65mL\\\beta=1.9GPa"

Put value formula

equation (1) put value


"1.9\times10^{9}=\frac{3.3\times10^6}{\frac{∆V}{65\times10^{-3}}}"

"1.9\times10^{9}=\frac{3.3\times10^6}{∆V}\times{{65\times10^{-3}}}"

"∆V=\frac{3.3\times10^6}{1.9\times10^{9}}\times{{65\times10^{-3}}}"

"∆V=1.128\times10^{-4}liter"